OpenStudy (narissa):

which number is between .5 and 8/9-1/3,25/36,22/24,7/24

OpenStudy (narissa):

is it 1/3

OpenStudy (narissa):

Can someone help me please!! I will give you a medal

OpenStudy (acxbox22):

calculate each and see which one works

OpenStudy (narissa):

I think its .33 1/3

OpenStudy (acxbox22):

is .33 between 0.5 and .89

OpenStudy (narissa):

yes

OpenStudy (narissa):

the other fractions into decimals are higher numbers

OpenStudy (acxbox22):

no, 0.33 is less than 0.5 so how can it be between 0.5 and 0.89

OpenStudy (acxbox22):

0.89=8/9

OpenStudy (narissa):

25/36=.694,22/24=.916 and 7/24=.2916 so I'm confused

OpenStudy (narissa):

did i do something wrong?

OpenStudy (narissa):

Can someone please help I'm confused

OpenStudy (narissa):

is it 25/36?

OpenStudy (narissa):

I will give someone a medal if they help me!

OpenStudy (texaschic101):

between 0.5 and 8/9 8/9 = 0.888 1/3 = 0.333 25/36 = 0.694 22/24 = 0.694 7/24 = 0.291 now which one is between 0.5 and 0.888 ?

OpenStudy (texaschic101):

oops...22/24 = 0.916

OpenStudy (narissa):

im confused i thought it was 1/3 but its not because thats smaller than 1/2 so then i thought it was 25/36 im just confused

OpenStudy (narissa):

i think it might be 22/24

OpenStudy (texaschic101):

25/36 = 0.694.....that is between 0.5 and 0.888 there are 2 ways to do this...either turn everything into fractions with a common denominator, then compare them...or turn everything into decimals, then compare them.

OpenStudy (narissa):

oh wow thank u can you help me with one more

OpenStudy (texaschic101):

sure...I can try

OpenStudy (narissa):

What is between 6/11 and .67 repeating

OpenStudy (narissa):

11/18,69/99,24/33,26/55

OpenStudy (narissa):

would it be 11/18 because that =.61

OpenStudy (texaschic101):

6/11 = 0.5454 so it has to be between 0.5454 and 0.6777 11/18 = 0.6111 69/99 = 0.6969 24/33 = 0.7272 26/55 = 0.4727 now look at your numbers...it has to be bigger then 0.5...so that eliminates the last one. It has to be less then 0.677...so that eliminates 2nd one and 3rd one...leaving us with 1st one as your answer....11/18

OpenStudy (narissa):

someone was helping me earlier and helped me gert the answers wrong

OpenStudy (texaschic101):

its 11/18....you cannot blame the person for helping you for giving the wrong answer...that person probably tried to really help. It is your own decision whether you want to use the answer you were given or try to figure it out with the help you are given.

OpenStudy (narissa):

ok thanks this is my first day on here

OpenStudy (texaschic101):

WELCOME :).....I think you will like it. Most of the people on here are helpful

OpenStudy (narissa):

do u know how I can find the quotient of 2 integers in simplest form of 4.62

OpenStudy (narissa):

I have 2 more problems to do

OpenStudy (narissa):

i think its 231/50 the original answer is 462/100 but have to simplify

OpenStudy (narissa):

does that sound right?

OpenStudy (texaschic101):

quotient of 2 integers is a division problem a/b = 4.62 a/b = 462/100 = 231/50

OpenStudy (narissa):

can u help me with 1 more?

OpenStudy (texaschic101):

what ya got ?

OpenStudy (narissa):

.36 repeating expressed as a quotient of 2 integers in simplest form

OpenStudy (narissa):

would it be 36/99 reduced is 4/11?

OpenStudy (narissa):

or 18/50=9/25

OpenStudy (texaschic101):

not too sure about this one because of the repeating decimal x = .3666 -- multiply by 100 100x = 36.666 x = .3666 -----------subtract 9x = 36 x = 36/9 which reduces to 4/1..... check this out.. http://www.ma.utexas.edu/users/mks/302/repdec.html

OpenStudy (narissa):

ok I got it its 4/11 thank u soooo much now I have to move on to more assignments on math thanks a whole bunch

OpenStudy (texaschic101):

YW :)

OpenStudy (narissa):

I have to get 8 assignments done today its not good

OpenStudy (texaschic101):

holy crap....you got your work cut out for you...good luck